# A compact model of Escherichia coli core and biosynthetic metabolism

### v2

@@ 3 Discussion
Here we presented iCH360, a medium-scale metabolic model of E. coli covering central and biosynthesis metabolism, together with the associated data and metabolic maps and results from several analysed use-cases. Similarly to previously constructed core models [orth_reconstruction_2010, hadicke_ecolicore2_2017], this model trades metabolic coverage for usability, interpretability, and ease of visualisation. It is well suited whenever a relatively small, highly curated network is desired, when computationally demanding analyses are to be performed, or as an educational tool in the field of metabolic modelling. When comparing some key properties of this model with those of its parent genome-scale model, we observed only small differences in the achievable biomass and product yields across a range of growth conditions, validating that, despite its contained size, the model captures the most salient metabolic features of the genome-scale network.
@@
To make this model easily usable in a variety of applications, we enriched the stoichiometric network structure by a curated layer of biological knowledge in the form of a graph data structure. This graph encodes information about biological entities in the network in a structured, ready-to-use format, including catalytic relationships between reactions and enzymes, the stoichiometric composition of protein complexes, and small-molecule regulation interactions. Further, we mapped to the model a range of quantitative parameters, including in vivo turnover number estimates and thermodynamic constants, extending the use of the model beyond simple stoichiometric analysis. A summary of the biological knowledge captured by iCH360 is shown in Table 3.
Owing to its medium size and the high level of curation, iCH360 lends itself to a wide range of modelling frameworks. Here, we demonstrated some representative examples. These include modelling the allocation of metabolic proteome via the enzyme-constraint variant of the model EC-iCH360, enumerating and analysing elementary flux modes in the network via the reduced model variant iCH360red, and performing thermodynamic-based analysis using the provided set of thermodynamic constants.
@@
While, at this stage, the parameterisation of the model is limited to turnover numbers and thermodynamic constants, we anticipate that additional parameter sets can easily be mapped to the model. Facilitated by extensive annotations present in the model and by the recent development of machine-learning-enabled kinetic constant estimators [kroll_deep_2021, li_deep_2022, kroll_turnover_2023, gollub_enkie_2023], a complete kinetic parameterisation of iCH360 is thus a valuable potential future development. In addition, probabilistic estimates of kinetic parameters [gollub_enkie_2023, lubitz_parameter_2019] can be combined with our existing thermodynamic parameterisation, making it possible to account for (potentially correlated) parameter uncertainty throughout the kinetic modelling process.
@@
In light of the above results, we believe that iCH360 holds the potential to become a reference metabolic model for E. coli.
@@ 4.1 Model assembly and curation
All the relevant pathways included in the model were assembled and curated based on information available in the EcoCyc [keseler_ecocyc_2017] and KEGG [kanehisa_kegg_2017] databases. The respective reactions were then extracted from iML1515 and parsed into a new model. In order to compute an equivalent biomass reaction, we first collected all the pathways required for the production of any component present in the iML1515 biomass reaction, but not in our model. These additional pathways (available in the repository supporting this manuscript) were manually curated based on available literature and database annotation to ensure they represent the biologically most relevant bioproduction route for each biomass component. By adding these pathways to iCH360, we obtained an extended model that was able to predict growth rates directly through the original iML1515 biomass reaction. The equivalent biomass function was then computed based on a reference flux distribution computed on this extended model, as explained in Supplementary Information A.2. Model assembly, manipulation and validation were performed using the COBRA Toolbox [ebrahim_cobrapy_2013]. Extension of database cross-annotation for the model reactions was performed through a mixture of automated database query and manual curation.
@@ 4.2 Network graphics
iCH360 can be visualised through a series of custom-built maps using the metabolic visualisation tool Escher [king_escher_2015] (see Figure 1). There are three main ways to visualise the model or solutions thereof. First, a complete map of the model, including all of its reactions, can be used (Figure 1). In order to provide a more compact representation of the network, a compressed second variant of the same map was constructed (Supplementary Figure 18). Here, long biosynthetic linear pathways were lumped into single pseudo-reactions, which only show the net production or consumption of metabolites by the pathway while omitting intermediates. Finally, individual maps for each of the main subsystems in the model are provided.
4.3 Graph data structure for linking reactions to enzymes and proteins
Information about the enzymes and proteins behind the stoichiometric model was collected in a knowledge graph. To build this graph, all available data on reaction-protein association and subunit composition were retrieved by automatic querying of the BioCyc database through its REST-based data retrieval API (https://biocyc.org/web-services.shtml). This information was then extended and curated based on a comparison with existing iML1515 GPR annotations. The resulting data were used to generate a directed graph wherein nodes represent biological entities (such as reactions, proteins, genes, and metabolites) and edges represent functional dependencies across them, including catalysis, subunit composition, post-translational modifications, and others. A complete list of node and edge types in the graph are provided in Supplementary Tables 1 and 2, respectively. All polypeptide nodes were annotated with their molecular mass (parsed from EcoCyc), enabling recursive computation of the molecular mass of any protein node in the graph (see Supplementary Information A.3). All manipulation and analyses of the graph data structure were performed using the NetworkX Python package [hagberg_exploring_2008], and the final data structure provided to the user in both Cytoscape and GML formats.
@@ 4.4 Primary and secondary catalytic edges
Catalytic edges in the graph, i.e. edges connecting reaction and enzyme nodes, were manually annotated as either primary or secondary, based on available evidence for the activity of each enzyme in the model with respect to its associated reactions. More specifically, a catalytic edge between a reaction-enzyme pair was labelled as secondary whenever the enzyme had been shown in literature to account for only minor catalytic activity for a reaction when compared to another isozyme. In this case, the references to the relevant literature were included as metadata for that edge. Whenever sufficient information was not available, all isozyme edges for a reaction were conservatively treated as primary.
@@ 4.5 Catalytic disruption analysis
For the catalytic disruption test, we identified condition-specific essential reactions in the following way: in the model, for each condition considered, we determined the reactions whose knockout led to inability to produce biomass. A small number of known false-positives, which are essential in iCH360 only due to its lack of certain reactions or pathways, were excluded from the analysis. For each essential reaction, we considered all associated genes and investigated the removal of each of them, individually, from the graph. The effect of the simulated knockout was propagated across the graph by removing all nodes from the graph for which the gene is required according to the nodes’ boolean GPR rules (see Supplementary Information A.3). Finally, the outcome of each simulated knockout was catalogued based on the reaction-level disruption it caused (see main text). Whenever multiple reactions were disrupted by the knockout of a gene, the strongest disruption among them was assigned to the knockout, in the following precedence order: complete disruption, full primary disruption, partial primary disruption, secondary disruption. The analysis was repeated for a total of $9$ growth conditions, and each condition-gene pair, labelled with the assigned disruption type, was compared with experimentally measured mutant relative fitness (averaged across available replicates for that condition-gene combination), using the dataset from [price_mutant_2018].
@@ 4.6 Construction of the enzyme-constrained metabolic model
In the sMOMENT formalism, positive and negative fluxes in each reaction are formally described, respectively, as positive fluxes in separate “forward” and “backward” versions of the reaction. To construct the enzyme-constrained model in sMOMENT format, reversible reactions in the model were therefore duplicated in the model to separately represent fluxes in forward or backward direction. Direction-specific turnover number estimates were parsed from [heckmann_kinetic_2020] and a default value of $65~{}\rm{s}^{-1}$ was used for transporters as in [heckmann_machine_2018]. To account for the fact that [heckmann_kinetic_2020] reports values as turnover numbers per polypeptide, the values were multiplied by the number of polypeptide subunits in each enzyme. For each reaction-enzyme pair, an enzyme cost per unit flux (in units of ${\rm g\cdot h\cdot mmol^{-1}}$) was then defined as:
@@
$\sum_{i}a_{i}~{}v_{i}\leq E_{\text{tot}}$ (2)
where $v_{i}$ denotes the flux in reaction $i$ (in ${\rm mmol~{}h^{-1}~{}gDW^{-1}}$) and $E_{\text{tot}}$ is a parameter denoting the total amount of enzyme mass (in g/gDW) that can be allocated to the flux mode. The constraint is enforced by augmenting the stoichiometric matrix of the model with an additional enzyme supply pseudoreaction (upper-bounded by $E_{\text{tot}}$) and an enzyme pool pseudometabolite consumed in each reaction with stoichiometry $e_{i}$ [bekiaris_automatic_2020].
@@ 4.7 Adjustment of turnover numbers across conditions
Turnover numbers from [heckmann_kinetic_2020] were adjusted based on the nonlinear programming (NLP) formulation detailed in Supplementary Information A.5). The reference flux distributions required by the procedure were obtained using the original (unadjusted) set of turnover numbers and the bounds on allowable adjustments ($\mathbf{l}$ and $\mathbf{u}$ in Eq. 35) set to ±$\pm 2$ (corresponding to a maximal 100-fold increase or reduction of each parameter from the original value). The linear program used to obtain reference flux distributions was formulated and solved with GUROBI [gurobi], while the nonlinear program used for turnover adjustment was formulated and solved with the open-source optimisation package CasADi [Andersson2019]. To investigate the effect of the ridge regularisation hyperparameter ($\rho$ in equation 35), we solved the adjustment problem for a broad range of values of this parameter and, each time, we computed the RMSE between measurements and predictions of enzyme abundance, as well as the mean absolute deviation of between original and adjusted turnover values, both computed for log10-transformed data (Supplementary Figure 11). Based on this information, a value of $\rho=1$, after which any further decrease in the amount of regularisation results in marginal reduction of the RMSE, was chosen to compute the final set of adjusted turnover numbers. Upon reparametrisation into apparent turnover numbers, (see Supplementary Information A.5.5), this set of adjusted parameters was used to parametrise the enzyme-constrained model variant EC-iCH360.
@@ 4.8 Enzyme allocation predictions
To validate enzyme allocation predictions against experimental values, we first retrieved measured polypeptide abundances for each growth condition from [schmidt_quantitative_2016] and imputed missing values using, whenever available, abundance values from the PAX Database [huang_paxdb_2023]. Then we estimated enzyme counts across conditions from polypeptide counts using non-negative least-squares estimation (Supplementary Information A.4). Next, we converted them into mass abundances (in units of ${\rm g~{}gDW^{-1}}$) based on the molecular mass of each complex (see Supplementary Information A.3) and assuming a cell dry mass of $2.8\times 10^{-13}\rm{g}$ (BIONUMBER ID 103904 [milo_bionumbers--database_2010]). Enzyme allocation predictions for each condition were then computed via EC-iCH360 by fixing the growth rate to the experimentally measured one from [schmidt_quantitative_2016] and minimising the total enzyme cost. To compute predictions using the turnover number estimates from [heckmann_kinetic_2020], the average saturation coefficient $\sigma$ was estimated, for each condition as:
@@
where $\mathcal{M}$ denotes the index set of model enzymes for which measurements are available and $\bar{E_{\rm{tot}}}$ is the total measured model enzyme abundance for that condition. This value of $\sigma$ was then used to scale predicted enzyme abundances before comparing them to the experimentally measured ones. To compute predictions with the adjusted parameter set, the scaling factors for each condition were obtained as part of the fitting procedure (see Supplementary Information A.5). All reported root mean squared errors were computed on log10 transformed enzyme abundances, excluding zero-predictions from the computation.

### v3

@@ 3 Discussion
Here we presented iCH360, a medium-scale metabolic model of E. coli covering central and biosynthesis metabolism, together with the associated data and metabolic maps and results from several analysed use cases. Similarly to previously constructed core models [orth_reconstruction_2010, hadicke_ecolicore2_2017], this model trades metabolic coverage for usability, interpretability, and ease of visualisation. It is well suited whenever a relatively small, highly curated network is desired, when computationally demanding analyses are to be performed, or as an educational tool in the field of metabolic modelling. When comparing some key properties of this model with those of its parent genome-scale model, we observed only small differences in the achievable biomass and product yields across a range of growth conditions, validating that, despite its contained size, the model captures the most salient metabolic features of the genome-scale network.
Further, we showed that the use of a well-curated, smaller-scale model can, in some instances, even correct unrealistic phenotypes predicted by its genome-scale parent. These unrealistic predictions from iML1515 are not the result of “errors” in the metabolic model. Rather, they are the result of applying simple stoichiometric methods, such as FBA, to a large network with many degrees of freedom. It is possible that the inclusion of additional constraints, such as thermodynamic feasibility under physiological conditions or proteome allocation bounds, would automatically render such solutions infeasible. However, these constraints, and the parameters required to implement them, are not always readily available. By assembling a smaller model and curating it with expert knowledge, we filter out many of these behaviours by construction, providing users with a versatile and interpretable tool to investigate central and biosynthetic metabolism in E. coli. Clearly, this comes at the cost of limited applicability in other scenarios, such as those where the metabolic subsystems not included in iCH360 (e.g. degradation pathways) are crucial to explaining or modelling a given phenotype. In these cases, the use of a genome-scale model (or an ad hoc reduction thereof) would still be an invaluable tool.
@@
To make this model easily usable in a variety of applications, we enriched the stoichiometric network structure with a curated layer of biological knowledge in the form of a knowledge graph. This graph encodes information about biological entities in the network in a structured, ready-to-use format, including catalytic relationships between reactions and enzymes, the stoichiometric composition of protein complexes, and small-molecule regulation interactions. In addition, we mapped to the model a range of quantitative parameters, including in vivo turnover number estimates and thermodynamic constants, extending the use of the model beyond a simple stoichiometric analysis. A summary of the biological knowledge captured by iCH360 is shown in Table 3.
Due to its medium size and the high level of curation, iCH360 lends itself to a wide range of modelling methods. Here, we demonstrated some representative examples. These include the calculation of production envelopes, the modelling of metabolic proteome allocation via the enzyme-constraint model variant EC-iCH360, enumerating and analysing elementary flux modes in the network via the reduced model variant iCH360red, and performing thermodynamic-based analysis using the set of thermodynamic constants provided. Nevertheless, we believe that many other analyses are possible. For example, alternative definitions of elementary pathways, such as elementary conversion modes, would be valuable to explore as a more tractable alternative to elementary flux modes [clement_unlocking_2021]. Similarly, our analyses of metabolic enzyme cost presented here have relied on simple, capacity-based definitions of enzyme cost, where a constant cost per unit flux is assumed for each reaction regardless of the condition studied. However, with additional kinetic parametrisation of the model, more complete enzyme cost functions, explicitly accounting for condition-specific metabolite concentrations and, consequently, enzyme costs, could be used to generate more realistic estimates of metabolic tradeoffs predicted by the model [noor_protein_2016].
@@
Indeed, while at this stage the parameterisation of the model is limited to turnover numbers and thermodynamic constants, we anticipate that additional parameter sets can easily be mapped to the model. Facilitated by extensive annotations present in the model and by the recent development of machine learning-enabled kinetic constant estimators [kroll_deep_2021, li_deep_2022, kroll_turnover_2023, gollub_enkie_2023], a complete kinetic parameterisation of iCH360 is thus a valuable potential future development. In addition, probabilistic estimates of the kinetic parameters [gollub_enkie_2023, lubitz_parameter_2019] can be combined with our existing thermodynamic parameterisation, making it possible to account for (potentially correlated) parameter uncertainty throughout the kinetic modelling process.
@@
In light of the above results, we believe that iCH360 has the potential to become a reference metabolic model for E. coli.
@@ 4.1 Model assembly and curation
All relevant pathways included in the model (the core metabolism reactions from ECC [orth_reconstruction_2010] and the biosynthesis pathways shown in Supplementary Figures S1-S4) were assembled and curated based on information available in the EcoCyc [keseler_ecocyc_2017] and KEGG [kanehisa_kegg_2017] databases. The respective reactions were then extracted from iML1515 and parsed into a new model. To compute an equivalent biomass reaction, we first collected all the pathways required for the production of the components present in the iML1515 biomass reaction (with the exception of a small number of compounds present with very small stoichiometry, which were neglected from this analysis for simplicity), but not in our model. We used the ”core” biomass reaction from iML1515 (BIOMASS_Ec_iML1515_core_75p37M), rather than alternative ”WT” reaction (BIOMASS_Ec_iML1515_WT_75p37M), since its smaller number of requirements made it easier to manually curate the pathways for their production. These additional pathways (available in the repository supporting this manuscript) were manually curated based on the available literature and database annotations to ensure they represent the biologically most relevant bioproduction route for each biomass component. By adding these pathways to iCH360, we obtained an extended model that was able to predict growth rates directly through the original iML1515 biomass reaction. The equivalent biomass function was then computed based on a reference flux distribution computed on this extended model, as explained in Supplementary Information A.2. Both growth-associated (the stoichiometry of intracellular ATP in the biomass reaction) and non-growth-associated (the lower bound on the maintenance reaction, ATPM) energy requirements were directly inherited from iML1515. Model assembly, manipulation and validation were performed using the COBRA Toolbox [ebrahim_cobrapy_2013]. The extension of database cross-annotation for the model reactions was performed through a mixture of automated database query and manual curation.
@@ 4.2 Network graphics
iCH360 can be visualised through a series of custom-built maps using the metabolic visualisation tool Escher [king_escher_2015] (see Figure 1). There are three main ways to visualise the model or solutions thereof. First, a complete map of the model, including all of its reactions, can be used (Figure 1). In order to provide a more compact representation of the network, a compressed second variant of the same map was constructed (Supplementary Figure S21). Here, long biosynthetic linear pathways were lumped into single pseudo-reactions, which only show the net production or consumption of metabolites by the pathway while omitting intermediates. Finally, individual maps are provided for each of the main subsystems in the model.
4.3 Production Envelope Analysis
All production envelopes shown in the main text and Supplementary Information were generated using COBRApy [ebrahim_cobrapy_2013]. For comparisons with ECC and ECC2, the maintenance requirement (lower bound on the ATPM reaction) of the two models was set to the same value used in iML1515 and iCH360 ($6.86$ mmol/gDW/h).
4.4 Knowledge graph for linking reactions to enzymes and proteins
Information about the enzymes and proteins behind the stoichiometric model was collected in a knowledge graph. To build this graph, all available data on reaction-protein association and subunit composition were retrieved by automatically querying the BioCyc database through its REST-based data retrieval API (https://biocyc.org/web-services.shtml). This information was then extended and curated based on a comparison with existing iML1515 GPR annotations. The resulting data were used to generate a directed graph in which nodes represent biological entities (such as reactions, proteins, genes, and metabolites) and edges represent functional dependencies across them, including catalysis, subunit composition, post-translational modifications, and others. A complete list of node and edge types in the graph is provided in Supplementary Tables 2 and 3, respectively. All polypeptide nodes were annotated with their molecular mass (parsed from EcoCyc), enabling recursive computation of the molecular mass of any protein node in the graph (see Supplementary Information A.3). All manipulation and analysis of the graph data structure were performed using the NetworkX Python package [hagberg_exploring_2008], and the final data structure is provided to the user in Cytoscape and GML formats.
@@ 4.5 Primary and secondary catalytic edges
Catalytic edges in the graph, i.e. edges connecting reaction and enzyme nodes, were manually annotated as either primary or secondary, based on available evidence for the activity of each enzyme in the model with respect to its associated reactions (Supplementary File S2). More specifically, a catalytic edge between a reaction-enzyme pair was labelled as secondary whenever the enzyme was shown in the literature to account for only minor catalytic activity for a reaction when compared with another isozyme. In this case, references to the relevant literature were included as metadata for that edge. Whenever sufficient information was not available, all isozyme edges for a reaction were conservatively treated as primary.
@@ 4.6 Catalytic disruption analysis
For the catalytic disruption test, we identified condition-specific essential reactions in the following way: in the model, for each condition considered, we determined the reactions whose knockout led to the inability to produce biomass. A small number of known false positives, which are essential in iCH360 only due to its lack of certain reactions or pathways, were excluded from the analysis. For each essential reaction, we considered all associated genes and investigated the removal of each of them individually from the graph. The effect of the simulated knockout was propagated across the graph by removing all nodes from the graph for which the gene is required according to the nodes’ Boolean GPR rules (see Supplementary Information A.3). Finally, the result of each simulated knockout was catalogued according to the reaction-level disruption it caused (see the main text). Whenever multiple reactions were disrupted by the knockout of a gene, the strongest disruption among them was assigned to the knockout, in the following precedence order: complete disruption, full primary disruption, partial primary disruption, secondary disruption. The analysis was repeated for a total of $9$ growth conditions, and each condition-gene pair, labelled with the assigned disruption type, was compared with experimentally measured mutant relative fitness (averaged across available replicates for that condition-gene combination), using the data set from [price_mutant_2018].
@@ 4.7 Construction of the enzyme-constrained metabolic model
In the sMOMENT formalism, the positive and negative fluxes in each reaction are formally described, respectively, as positive fluxes in separate “forward” and “backward” versions of the reaction. To construct the enzyme-constrained model in the sMOMENT format, reversible reactions in the model were duplicated in the model to separately represent fluxes in forward or backward direction. Direction-specific turnover number estimates were parsed from [heckmann_kinetic_2020] and a default value of $65~{}\rm{s}^{-1}$ was used for transporters as in [heckmann_machine_2018]. To account for the fact that [heckmann_kinetic_2020] reports values as turnover numbers per polypeptide, the values were multiplied by the number of polypeptide subunits in each enzyme. For each reaction-enzyme pair, an enzyme cost per unit flux (in units of ${\rm g\cdot h\cdot mmol^{-1}}$) was then defined as:
@@
$\sum_{i}a_{i}~{}v_{i}\leq e_{\rm tot}$ (2)
where $v_{i}$ denotes the flux in reaction $i$ (in ${\rm mmol~{}h^{-1}~{}gDW^{-1}}$) and $e_{\rm tot}$ is a parameter denoting the total amount of enzyme mass (in g/gDW) that can be allocated to the flux mode. The constraint is enforced by augmenting the stoichiometric matrix of the model with an additional enzyme supply pseudoreaction (upper-bounded by $e_{\rm tot}$) and an enzyme pool pseudometabolite consumed in each reaction with stoichiometry $a_{i}$ [bekiaris_automatic_2020].
@@ 4.8 Adjustment of turnover numbers across conditions
Turnover numbers from [heckmann_kinetic_2020] were adjusted based on the nonlinear programming (NLP) formulation detailed in Supplementary Information A.5). The reference flux distributions required by the procedure were obtained using the original (unadjusted) set of turnover numbers and the bounds on allowable adjustments ($\mathbf{u_{\rm min}}$ and $\mathbf{u_{\rm max}}$ in Eq. 34) set to ±$\pm 2$ (corresponding to a maximal 100-fold increase or reduction of each parameter from the original value). The linear program used to obtain reference flux distributions was formulated and solved with GUROBI [gurobi], while the nonlinear program used for turnover adjustment was formulated and solved with the open-source optimisation package CasADi [Andersson2019]. To investigate the effect of the ridge regularisation hyperparameter ($\rho$ in equation (34)), we solved the adjustment problem for a broad range of values of this parameter and, each time, we computed the RMSE between measurements and predictions of enzyme abundance, as well as the mean absolute deviation of between original and adjusted turnover values, both computed for log10-transformed data (Supplementary Figure S13). Based on this information, a value of $\rho=1$, after which any further decrease in the amount of regularisation results in marginal reduction of the RMSE, was heuristically chosen to compute the final set of adjusted turnover numbers. Upon reparametrisation into apparent turnover numbers, (see Supplementary Information A.5.5), this set of adjusted parameters was used to parametrise the enzyme-constrained model variant EC-iCH360. For the leave-one-out cross validation, the adjustment procedure was run multiple times, each time excluding abundance data from a condition from the training dataset. The resulting adjusted parameter set was then used to generate enzyme abundance predictions for the condition left out, which were then compared against measured values.
@@ 4.9 Enzyme allocation predictions
To validate enzyme allocation predictions against experimental values, we first retrieved measured polypeptide abundances for each growth condition from [schmidt_quantitative_2016] and imputed missing values using, whenever available, abundance values from the PAX Database [huang_paxdb_2023]. Then we estimated enzyme counts across conditions from polypeptide counts using non-negative least-squares estimation (Supplementary Information A.4). Next, we converted them into mass abundances (in units of ${\rm g~{}gDW^{-1}}$) based on the molecular mass of each complex (see Supplementary Information A.3) and assuming a cell dry mass of $2.8\times 10^{-13}\rm{g}$ (BIONUMBER ID 103904 [milo_bionumbers--database_2010]). Enzyme allocation predictions for each condition were then computed via EC-iCH360 by fixing the growth rate to the experimentally measured one from [schmidt_quantitative_2016] and minimising the total enzyme cost, initially using a value of $1$ for the average saturation parameter $\sigma$ (Eq. (1)). For the predictions computed using the turnover number estimates from [heckmann_kinetic_2020], the average saturation coefficient $\sigma$ was then estimated from data, for each condition, as:
@@
where $\mathcal{M}$ denotes the index set of model enzymes for which measurements are available, $e_{i}$ is the predicted abundance for the $i$th enzyme, and $\bar{e}_{\rm{tot}}$ is the total measured model enzyme abundance for that condition. This value of $\sigma$ was then used to scale the predicted enzyme abundances before comparing them with the ones measured experimentally. Note that this choice of $\sigma$ ensures that the sum of predicted abundances matches that of measured ones. To compute predictions with the adjusted parameter set, the scaling factors for each condition were obtained as part of the fitting procedure (see Supplementary Information A.5). All reported root mean squared errors were computed on log10 transformed enzyme abundances, excluding enzymes with zero predicted abundance from the computation.